OpenStudy (anonymous):

Find a Cartesian equation relating x and y corresponding to the parametric equations x=5sin(5t)y=9cos(5t) Write your answer in the form P(xy)=0 where P(xy) is a polynomial in x and y such that the coefficient of y^2 is 25.

OpenStudy (anonymous):

Hint: It's an ellipse.

OpenStudy (anonymous):

I found that when I looked at the graph but the part that confuses me is when it says the coefficient on y^2 is 25

OpenStudy (anonymous):

So you have\[ ax^2+25y^2=c \]

OpenStudy (anonymous):

Meaning you probably want \(a=9^2=81\)

OpenStudy (anonymous):

That leaves you with \(c=(25)(81)\)

OpenStudy (anonymous):

Ok thanks that helped with the question.